Island Novelties, Inc., of Palau makes two products—Hawaiian Fantasy and Tahitian Joy. Each product's selling price, variable expense per unit and annual sales volume are as follows:

Hawaiian Fantasy Tahitian Joy
Selling price per unit $ 30 $ 125
Variable expense per unit $ 21 $ 25
Number of units sold annually 10,000 5,600

Fixed expenses total $565,500 per year.

Required:

1. Assuming the sales mix given above, do the following:

a. Prepare a contribution format income statement showing both dollar and percent columns for each product and for the company as a whole.

b. Compute the company's break-even point in dollar sales. Also, compute its margin of safety in dollars and its margin of safety percentage.

2. The company has developed a new product called Samoan Delight that sells for $50 each and that has variable expenses of $35 per unit. If the company can sell 20,000 units of Samoan Delight without incurring any additional fixed expenses:

a. Prepare a revised contribution format income statement that includes Samoan Delight. Assume that sales of the other two products does not change.

b. Compute the company’s revised break-even point in dollar sales. Also, compute its revised margin of safety in dollars and margin of safety percentage.
Answer #1

Ans.

1(a)

Contribution format income statement showing both dollar and percent columns for each product and for the company as a whole.

Hawaiian Fantasy Tahitian Joy Total (Company )
Amount ($) % Amount ($) % Amount ($) %
Sales (a) $ 300,000.00 100% $ 700,000.00 100% $ 1,000,000.00 100%
Variable Expenses (b) $ 210,000.00 70% $ 140,000.00 20% $      350,000.00 35%
Contribution Margin(c) = (a-b) $    90,000.00 30% $ 560,000.00 80% $      650,000.00 65%
Fixed Expenses (d) $      565,500.00
Net Operating Income (e)=(c-d) $        84,500.00

1(b)

Break even point (in Dollar Sales)

=Fixed expenses / Contribution Margin Ratio = $565,500 / 0.65 = $ 870,000

Margin of safety (in dollars) = Actual sales – Break-even sales

= $ 1,000,000 - $ 870,000 = $ 130,000

Margin of safety percentage = Margin of safety in dollars / Actual sales

= $130,000/$ 1,000,000 = 13%

2(a)

Revised contribution format income statement that includes Samoan Delight.

Hawaiian Fantasy Tahitian Joy Samoan Delight Total (Company )
Amount ($) % Amount ($) % Amount ($) % Amount ($) %
Sales (a) $ 300,000.00 100% $ 700,000.00 100% $ 1,000,000.00 100% $ 2,000,000.00 100%
Variable Expenses (b) $ 210,000.00 70% $ 140,000.00 20% $      700,000.00 70% $ 1,050,000.00 52.5%
Contribution Margin(c) = (a-b) $    90,000.00 30% $ 560,000.00 80% $      300,000.00 30% $      950,000.00 47.5%
Fixed Expenses (d) $      565,500.00
Net Operating Income (e)=(c-d) $      384,500.00

2(b)

Break-even point ( in dollar sales )

Fixed expenses / Contribution Margin Ratio = $565,500 / 0.475 = $ 1,190,526.32

Margin of safety (in dollars) = Actual sales – Break-even sales

= $ 2,000,000 - $1,190,526.32 = $ 809,473.68

Margin of safety percentage = Margin of safety in dollars / Actual sales

= $809,473.68 /$ 2,000,000 =40.47%

Note = Sales is calculated as No. of units sold * Selling price per unit

And Variable Cost is calculated as No. of units sold * Variable Expense per unit